# Active Causal Structure Learning with Latent Variables: Towards Learning to Detour in Autonomous Robots

### v1

@@ Abstract
Keywords:
Artificial General Intelligence Causal Discovery Theory of Surprise Dynamic Decision Network MEU POMDP Causal Artificial Intelligence.
@@ 2.1 Partially Observed Markov Decision Process (POMDP)
  • $A$ is the set of actions that the agent can take.
@@ 2.3.1 Causal relations between random variables
$\left\{\begin{array}[]{l}R=X\xrightarrow{}Y\Rightarrow\exists\;f\wedge\exists\;\mathcal{N}_{y}\;\;|\;Y:=f(X,\mathcal{N}_{y})\\ \\ R=X\leftarrow{}Y\Rightarrow\exists\;f\wedge\exists\;\mathcal{N}_{x}\;\;|\;X:=f(Y,\mathcal{N}_{x})\\ \\ R=X\leftrightarrow Y\Rightarrow\exists\;f,g\wedge\exists\;\mathcal{N}_{x},\mathcal{N}_{y},Z\;\;|\;X:=f(Z,\mathcal{N}_{x})\wedge Y:=g(Z,\mathcal{N}_{y})\\ \\ R=X\perp Y\Rightarrow X\text{ and }Y\text{are independent, there is no function that relates them to each other }\par\end{array}\right.$
@@ 2.3.2 Causal relations between actions and random variables
$\left\{\begin{array}[]{l}R=D\xrightarrow{}O\Rightarrow\text{The activation function of $D$ influences the values that $O$ takes}\\ \\ R=D\perp O\Rightarrow\text{The activation function of $D$ does not influence the values that $O$ takes}\par\end{array}\right.$
@@ 3.2 Environment description
  • $Target\;position$: $(x,y)$ coordinates of the Target. They are set to be $(x_{T},y_{T})=(10,7.5)$ always.
  • $Agent\;position$: $(x,y)$ coordinates of the center of the Agent, initially they are $(x_{A},y_{A})=(1,7.5)$.
  • $Agent\;orientation$: Orientation where the Agent is looking at. In this paper, it will be 00 always.
  • $Agent\;Width$ and $Agent\;Shape$: These variables indicate the shape and width of the agent. A square of side $0.75$ has been chosen for the agent.
  • $Barrier\;exist$: Boolean variable indicating whether or not the environment has a barrier.
  • $Start\;Barrier$ and $End\;Barrier$ coordinates: Coordinates of the beginning and end of the barrier. The barrier will consist of a segment between these points, made of spikes. The coordinates are $Start_{B}=(4.5,1.5)$ and $End_{B}=(4.5,15)$.
  • $Spikes\;separation$ and $spikes\;length$: Separation of the barrier’s spikes and their length. We set $Spikes\;separation=0.5$ to prevent the agent from passing through the barrier due to its width, and $spikes\;length=0.5$ defines the length of each spike.
@@
$Target\;position$ $(10,7.5)$
$Agent\;position$ $(1,7.5)$
$Agent\;orientation$ 00
$Agent\;width$ $0.75$
$Agent\;shape$ square
$Barrier\;Exist$ Yes
$Start\;Barrier$ $(4.5,1.5)$
$End\;Barrier$ $(4.5,15)$
$Spikes\;separation$ $0.5$
$Spikes\;length$ $0.5$
@@ 3.3 Actions and Observations Space
where $(x,y)$ are the coordinates $Agent\;position$, $\alpha$ is the value of $Agent\;orientation$ and $s$ the number of steps taken by the agent.
$Step\;Aside$:
This action allows the agent to move perpendicular to the direction in which it is facing. The decision domain is $[-2.5,2.5]$, negative values are for moving to the left and positive values for moving to the right. The mechanistic function behind it is the following:
$\begin{split}Step\;Aside&\colon\mathbb{R}^{2}\times[-\pi,\pi]\times[-2.5,2.5]\longrightarrow\mathbb{R}^{2}\\ SA(x,y,\alpha,s)=&\left\{\begin{array}[]{ll}(x,y)+s(\cos{(\alpha}-\frac{\pi}{2}),\sin{(\alpha-\frac{\pi}{2})})&\text{si}\;s<0\\ \\ (x,y)+s(\cos{(\alpha+\frac{\pi}{2})},\sin{(\alpha+\frac{\pi}{2}}))&\text{si}\;s\geq 0\end{array}\right.\end{split}$ (5)
@@
  • $Depth$: It corresponds to the Euclidean distance between the agent and the target.
@@
  • $Heading\;angle$: This variable measures the angle between the agent and the target, using the agent coordinates as the reference system.
@@
  • $Barrier\;Tactile$: This is a binary variable that is $1$ if the infinity norm between the center of the agent and any of the spikes of the barrier is lower than $2$, and 00 otherwise.
    $BT=\left\{\begin{array}[]{lc}1&\text{if}\;\underset{s\in spikes}{\max}||(x_{A},y_{A})-(x_{s},y_{s})||_{\infty}\leq 2\\ \\ 0&\text{otherwise }\end{array}\right.$ (8)
  • $Target\;in\;visual\;Field$: A binary variable, it will be $1$ if $Depth$ is lower than $2$ and $Heading\;angle$ is between $[-\frac{\pi}{2},\frac{\pi}{2}]$.
@@ 3.4 Agent´s initial Dynamic Decision Network
where $sf$ and $sa$ refer to the decision values that the actions $Step\;Forward$ and $Step\;Aside$ take, respectively.
(a) $Step\_Forward$ energy function (11)
(b) $Step\_Aside$ energy function (12)
@@ 3.5 Problem modeled as a POMDP
  • $\Omega=\{Depth\;,Heading\;Angle\;,Barrier\;Tactile,\;Target$ $\;in\;Visual\;Field\}$.
@@
  • Since we are going to use a DDN that contemplates only the next instant in the future, the $\gamma$ factor is equal to 00.
@@ 4.2.2 Mathematical formalism of surprise divergence measure
    $H(Q)-H(P)$ measures the difference in the uncertainty given by both distributions.
@@
$D_{KL}(Q||P)\xrightarrow{}0$Leftrightarrow $D_{S}(Q||P)^{2}\xrightarrow{}0$.
@@ 5.1.1 Utility Surprise
First of all, recall that the agent’s way of acting is summarized by observing the values of the observation variables $\textbf{x}_{t-1}$ at time $t-1$ and computing the probabilities on the observation variables at time $t$ from these values, as well as the possible values of the action space $Act_{t-1}$, denoted by $\text{I\kern-1.49994ptP}(\textbf{x}_{t}|\textbf{x}_{t-1},Act_{t-1})$. This enables the agent to calculate all possible values of $U(\textbf{x}_{t})$ and then choose the decision $\textbf{a}^{*}$ that brings the maximum expected utility,
@@
The way the agent will detect the latent variables will be by comparing the utility value it receives at time $t$ with the probability distribution $P(\textbf{x}_{t}|\textbf{x}_{t-1},Act_{t-1})$ it computed at time $t-1$. Two critical aspects must be taken into account in this comparison:
@@
As for the case $2$, if the utility obtained is higher than expected, then we would have the opposite case, i.e. it has been discovered that there is a variable that influences and helps to obtain higher utility values, so it is convenient to learn to model when this may occur. However, in this paper, we will focus on detecting the latent variables that have a negative influence on the utility, and therefore the probability of influence will be close to 00, in the case that the utility obtained is higher than expected.
If the surprise coefficient is close to 00, as for the cases $3$ and $4$, then it means that the probability that there is a variable that the agent does not observe, and that changes its interaction with the environment will be low, since the agent is able to correctly predict the values of the observation variables that it will get after executing the actions. Taking all this into account, the coefficient that will be responsible for making this comparison will be an extrapolation of the one described in (22), which will also include whether the utility obtained is higher or lower than the one expected.
@@ 5.1.2 Probability of Influence of the Hidden Variable
The probability of influence of the latent variable will be a function dependent on the value of the coefficient of the utility surprise. The range of possible values that $C_{U}$ can take is potentially (,)($\infty$,$\infty$)( - ∞ , ∞ ), and the goal is to convert it to the range $[0,1]$. Therefore, assuming we want to calculate $\text{I\kern-1.49994ptP}(HV=0)$, the lower the coefficient, the greater the negative surprise, that is, the greater the evidence that $HV=1$, so $\text{I\kern-1.49994ptP}(HV=0)$ must tend to 00 . On the other hand, the higher it is, the greater the evidence that $HV=0$, thus $\text{I\kern-1.49994ptP}(HV=0)$ must tend to $1$. However, when $C_{U}$ is close to 00, this will indicate that there has not been enough surprise at instant t, so we do not have enough evidence to say that $HV=1$, but neither is it correct to say that $HV=0$, since it could happen that the influence of the latent variable is practically null, therefore, when $C_{U}=0$ we will have that $\text{I\kern-1.49994ptP}(HV=0)=\frac{1}{2}$.
Summing up, the range of possible values that $C_{U}$ can take is potentially (,)($\infty$,$\infty$)( - ∞ , ∞ ), and the goal is to convert it to the range $[0,1]$ as long as $\text{I\kern-1.49994ptP}(HV=0|C_{U}=0)=\frac{1}{2}$, $\text{I\kern-1.49994ptP}(HV=0|C_{U}\xrightarrow{}\infty)=1$ and $\text{I\kern-1.49994ptP}(HV=0|C_{U}\xrightarrow{}-\infty)=0$. This can be achieved by different functions (such as the logistic, hyperbolic tangent, normal cumulative distribution function, etc). However, we will choose a function that is consistent with the underlying statistical theory behind the coefficient. The absolute value $|C_{U}|$, based on the theorem 4.2, would statistically follow a $\chi$ distribution with one degree of freedom, since its the absolute value of something that tend to a normal distribution. Therefore, we will use the cumulative distribution function of a $\chi$ distribution as a baseline. In addition, we must consider the sign of the coefficient and guarantee that $C_{U}=0$ corresponds to a probability of $\frac{1}{2}$. To this end, we have developed the following function:
@@
which behaves as we have previously described. That is, when $C_{U}<<0$, the probability of no influence will tend to 00; meanwhile if $C_{U}>>0$, the probability will tend to $1$.
@@ 5.1.3 Selection of related variables: Surprise in Observations
$\;\;\;\;\;H_{0}:\text{The variable }Obs\text{ has not experienced a surprise in time }t$
$H_{1}:\text{The variable }Obs\text{ has experienced a surprise in time }t,$
@@
Returning to the Learning to Detour case (Figure 5), when the agent first bumps into the barrier, an unexpected effect on the $Barrier\;Tactile$ Observation variable occurs, since it is not predicted by any existing transition. Similarly, an unexpected effect on the $Depth$ Observation variable occurs after the activation of the $Step\;Forward$ Action variable, since the associated transition predicts a reduction on the $Depth$ value, yet the $Depth$ value remains unaltered since no forward movement can take place, due to the blocking effect by the barrier. For the purposes of visualization, the surprises were normalized by dividing them by the sum of the surprises present in the observed variables. This allowed for the accurate representation (Figure 5) of the fact that the variables $Barrier\;Tactile$ and $Depth$ experienced the greatest degree of surprise, and therefore represent the variables that should be selected.
@@
(a) Surprises $Barrier\;Tactile$
(b) Surprises $Depth$
(c) Surprises $Heading\;angle$
(d) Surprises $Target\;in\;Visual\;Field$
@@ 5.2 CPT Estimation
Compute Probability of Influence of $HV$ ($IP_{HV}$) for each observation $i$ in dataset $D$ ;
@@
Expectation 00: Build dataset $D^{\prime}_{0}$ from $D$, including imputed values of $HV$ using $IP_{HV}$ and weights $w_{i}$
Maximization 00 : Learn parameters $\Theta_{0}$ of $DDN_{0}$ by MLE or MAP;
@@
       Expectation-step: Generate new dataset $D^{\prime}_{t}$ with imputed values for $HV$ using $DDN_{t-1}$;
@@ 5.3 Learning to Detour behavior
Figure 8: Trajectories before and after learning $HV$
The first thing to notice, is the shift from the agent’s exclusive reliance on the $Step\;Forward$ action (before learning) to a more nuanced approach (during learning), involving a reduction in the power of $Step\;Forward$ decisions and the initiation of $Step\;Aside$ actions to the right. This is due to the changes made to the CPT of $BT_{t+1}$. As can be seen in the marginalized CPT of $BT_{t+1}$ before and after learning (Table 2), there is a greater probability of a $BT=1$ in the presence of the $HV$. This CPT, along with the CPT learned from $HV$ (Table 3), enables the agent to act in this way.
@@
The second significant event occurs when the agent collides with the barrier. During learning, the agent reduces the power of the $Step\;Forward$ action. This is a consequence of the agent’s adaptation, which has led to a reduction in the expected utility of activating $Step\;Forward$ when the probability of $HV=1$ is high. It is important to notice that at this moment, another important change takes place within our framework. That is, one of the key objectives is to enable the agent to learn from unexpected outcomes, which will enable it to make more accurate predictions and, therefore, avoid surprises in the future. In our case, Figures 9 and 10 illustrate that the surprises associated with $BT$ (Barrier Tactile) and $D$ (Depth) are reduced in comparison to those observed prior to learning. This clearly indicates that the predictions made by the agent are more accurate with respect to the environment when the $HV$ factor is present. In summary, before learning, the agent tends to go straight to the barrier and bumps into it in many occasions. Yet, after learning, the agent tends to avoid the barrier by side stepping around it in its trajectory to the target.
Figure 9: Surprise coefficients over $BT$ (Barrier Tactile) before and after learning $HV$.
Figure 10: Surprise coefficients over $D$ (Depth) before and after learning $HV$.
Figure 11: Surprise coefficients over the Utility before and after learning $HV$.

### v2

# Working Paper: Active Causal Structure Learning with Latent Variables: Towards Learning to Detour in Autonomous Robots

@@ 2.1 Partially Observed Markov Decision Process (POMDP)
  • AA is the set of actions that the agent can take.
@@ 2.3.1 Causal relations between random variables
[…]
@@ 2.3.2 Causal relations between actions and random variables
[…]
@@ 3.2 Environment description
  • TargetpositionTarget\;position: $(x,y)$ coordinates of the Target. They are set to be $(x_{T},y_{T})=(10,7.5)$ always.
  • AgentpositionAgent\;position: $(x,y)$ coordinates of the center of the Agent, initially they are $(x_{A},y_{A})=(1,7.5)$.
  • AgentorientationAgent\;orientation: Orientation where the Agent is looking at. In this paper, it will be 0 always.
  • AgentWidthAgent\;Width and AgentShapeAgent\;Shape: These variables indicate the shape and width of the agent. A square of side $0.75$ has been chosen for the agent.
  • BarrierexistBarrier\;exist: Boolean variable indicating whether or not the environment has a barrier.
  • StartBarrierStart\;Barrier and EndBarrierEnd\;Barrier coordinates: Coordinates of the beginning and end of the barrier. The barrier will consist of a segment between these points, made of spikes. The coordinates are $Start_{B}=(4.5,1.5)$ and $End_{B}=(4.5,15)$.
  • SpikesseparationSpikes\;separation and spikeslengthspikes\;length: Separation of the barrier’s spikes and their length. We set Spikesseparation=0.5Spikes\;separation=0.5 to prevent the agent from passing through the barrier due to its width, and spikeslength=0.5spikes\;length=0.5 defines the length of each spike.
@@
TargetpositionTarget\;position $(10,7.5)$
AgentpositionAgent\;position $(1,7.5)$
AgentorientationAgent\;orientation 0
AgentwidthAgent\;width $0.75$
AgentshapeAgent\;shape square
BarrierExistBarrier\;Exist Yes
StartBarrierStart\;Barrier $(4.5,1.5)$
EndBarrierEnd\;Barrier $(4.5,15)$
SpikesseparationSpikes\;separation $0.5$
SpikeslengthSpikes\;length $0.5$
@@ 3.3 Actions and Observations Space
where $(x,y)$ are the coordinates AgentpositionAgent\;position, $\alpha$ is the value of AgentorientationAgent\;orientation and $s$ the number of steps taken by the agent.
StepAsideStep\;Aside:
This action allows the agent to move perpendicular to the direction in which it is facing. The decision domain is [2.5,2.5][-2.5,2.5], negative values are for moving to the left and positive values for moving to the right. The mechanistic function behind it is the following:
[…]
@@
  • DepthDepth: It corresponds to the Euclidean distance between the agent and the target.
@@
  • HeadingangleHeading\;angle: This variable measures the angle between the agent and the target, using the agent coordinates as the reference system.
@@
  • BarrierTactileBarrier\;Tactile: This is a binary variable that is $1$ if the infinity norm between the center of the agent and any of the spikes of the barrier is lower than $2$, and 0 otherwise.
[…]
  • TargetinvisualFieldTarget\;in\;visual\;Field: A binary variable, it will be $1$ if $Depth$ is lower than $2$ and HeadingangleHeading\;angle is between $\frac{\pi}{2},\frac{\pi}{2}]$.
@@ 3.4 Agent´s initial Dynamic Decision Network
where $sf$ and $sa$ refer to the decision values that the actions StepForwardStep\;Forward and StepAsideStep\;Aside take, respectively.
(a) Step_ForwardStep\_Forward energy function (11)
(b) Step_AsideStep\_Aside energy function (12)
@@ 3.5 Problem modeled as a POMDP
  • $\Omega=\{Depth\;,Heading\;Angle\;,Barrier\;Tactile,\;Target$ inVisualField}\;in\;Visual\;Field\}.
@@
  • Since we are going to use a DDN that contemplates only the next instant in the future, the $\gamma$ factor is equal to 0.
@@ 4.2.2 Mathematical formalism of surprise divergence measure
    H(Q)H(P)H(Q)-H(P) measures the difference in the uncertainty given by both distributions.
@@
$D_{KL}(Q||P)\xrightarrow{}0$ $\Leftrightarrow$ $D_{S}(Q||P)^{2}\xrightarrow{}0$.
@@ 5.1.1 Utility Surprise
First of all, recall that the agent’s way of acting is summarized by observing the values of the observation variables $\textbf{x}_{t-1}$ at time t1t-1 and computing the probabilities on the observation variables at time $t$ from these values, as well as the possible values of the action space $Act_{t-1}$, denoted by $\text{I\kern-1.49994ptP}(\textbf{x}_{t}|\textbf{x}_{t-1},Act_{t-1})$. This enables the agent to calculate all possible values of $U(\textbf{x}_{t})$ and then choose the decision $\textbf{a}^{*}$ that brings the maximum expected utility,
@@
The way the agent will detect the latent variables will be by comparing the utility value it receives at time $t$ with the probability distribution $P(\textbf{x}_{t}|\textbf{x}_{t-1},Act_{t-1})$ it computed at time t1t-1. Two critical aspects must be taken into account in this comparison:
@@
As for the case $2$, if the utility obtained is higher than expected, then we would have the opposite case, i.e. it has been discovered that there is a variable that influences and helps to obtain higher utility values, so it is convenient to learn to model when this may occur. However, in this paper, we will focus on detecting the latent variables that have a negative influence on the utility, and therefore the probability of influence will be close to 0, in the case that the utility obtained is higher than expected.
If the surprise coefficient is close to 0, as for the cases $3$ and $4$, then it means that the probability that there is a variable that the agent does not observe, and that changes its interaction with the environment will be low, since the agent is able to correctly predict the values of the observation variables that it will get after executing the actions. Taking all this into account, the coefficient that will be responsible for making this comparison will be an extrapolation of the one described in (22), which will also include whether the utility obtained is higher or lower than the one expected.
@@ 5.1.2 Probability of Influence of the Hidden Variable
The probability of influence of the latent variable will be a function dependent on the value of the coefficient of the utility surprise. The range of possible values that $C_{U}$ can take is potentially $(-\infty,\infty)$, and the goal is to convert it to the range $[0,1]$. Therefore, assuming we want to calculate $\text{I\kern-1.49994ptP}(HV=0)$, the lower the coefficient, the greater the negative surprise, that is, the greater the evidence that $HV=1$, so $\text{I\kern-1.49994ptP}(HV=0)$ must tend to 0 . On the other hand, the higher it is, the greater the evidence that $HV=0$, thus $\text{I\kern-1.49994ptP}(HV=0)$ must tend to $1$. However, when $C_{U}$ is close to 0, this will indicate that there has not been enough surprise at instant t, so we do not have enough evidence to say that $HV=1$, but neither is it correct to say that $HV=0$, since it could happen that the influence of the latent variable is practically null, therefore, when $C_{U}=0$ we will have that $\text{I\kern-1.49994ptP}(HV=0)=\frac{1}{2}$.
Summing up, the range of possible values that $C_{U}$ can take is potentially $(-\infty,\infty)$, and the goal is to convert it to the range $[0,1]$ as long as $\text{I\kern-1.49994ptP}(HV=0|C_{U}=0)=\frac{1}{2}$, $\text{I\kern-1.49994ptP}(HV=0|C_{U}\xrightarrow{}\infty)=1$ and $\text{I\kern-1.49994ptP}(HV=0|C_{U}\xrightarrow{}-\infty)=0$. This can be achieved by different functions (such as the logistic, hyperbolic tangent, normal cumulative distribution function, etc). However, we will choose a function that is consistent with the underlying statistical theory behind the coefficient. The absolute value $|C_{U}|$, based on the theorem 4.2, would statistically follow a $\chi$ distribution with one degree of freedom, since its the absolute value of something that tend to a normal distribution. Therefore, we will use the cumulative distribution function of a $\chi$ distribution as a baseline. In addition, we must consider the sign of the coefficient and guarantee that $C_{U}=0$ corresponds to a probability of $\frac{1}{2}$. To this end, we have developed the following function:
@@
which behaves as we have previously described. That is, when $C_{U}<<0$, the probability of no influence will tend to 0; meanwhile if $C_{U}>>0$, the probability will tend to $1$.
@@ 5.1.3 Selection of related variables: Surprise in Observations
[…]
@@
Returning to the Learning to Detour case (Figure 5), when the agent first bumps into the barrier, an unexpected effect on the BarrierTactileBarrier\;Tactile Observation variable occurs, since it is not predicted by any existing transition. Similarly, an unexpected effect on the $Depth$ Observation variable occurs after the activation of the StepForwardStep\;Forward Action variable, since the associated transition predicts a reduction on the $Depth$ value, yet the $Depth$ value remains unaltered since no forward movement can take place, due to the blocking effect by the barrier. For the purposes of visualization, the surprises were normalized by dividing them by the sum of the surprises present in the observed variables. This allowed for the accurate representation (Figure 5) of the fact that the variables BarrierTactileBarrier\;Tactile and $Depth$ experienced the greatest degree of surprise, and therefore represent the variables that should be selected.
@@
(a) Surprises BarrierTactileBarrier\;Tactile
(b) Surprises DepthDepth
(c) Surprises HeadingangleHeading\;angle
(d) Surprises TargetinVisualFieldTarget\;in\;Visual\;Field
@@ 5.2 CPT Estimation
Compute Probability of Influence of HVHV ($IP_{HV}$) for each observation $i$ in dataset $D$ ;
@@
Expectation 0: Build dataset $D^{\prime}_{0}$ from $D$, including imputed values of HVHV using $IP_{HV}$ and weights $w_{i}$
Maximization 0 : Learn parameters $\Theta_{0}$ of $DDN_{0}$ by MLE or MAP;
@@
 Expectation-step: Generate new dataset $D^{\prime}_{t}$ with imputed values for HVHV using $DDN_{t-1}$;
@@ 5.3 Learning to Detour behavior
Figure 8: Trajectories before and after learning HVHV
The first thing to notice, is the shift from the agent’s exclusive reliance on the StepForwardStep\;Forward action (before learning) to a more nuanced approach (during learning), involving a reduction in the power of StepForwardStep\;Forward decisions and the initiation of StepAsideStep\;Aside actions to the right. This is due to the changes made to the CPT of $BT_{t+1}$. As can be seen in the marginalized CPT of $BT_{t+1}$ before and after learning (Table 2), there is a greater probability of a $BT=1$ in the presence of the $HV$. This CPT, along with the CPT learned from $HV$ (Table 3), enables the agent to act in this way.
@@
The second significant event occurs when the agent collides with the barrier. During learning, the agent reduces the power of the StepForwardStep\;Forward action. This is a consequence of the agent’s adaptation, which has led to a reduction in the expected utility of activating StepForwardStep\;Forward when the probability of $HV=1$ is high. It is important to notice that at this moment, another important change takes place within our framework. That is, one of the key objectives is to enable the agent to learn from unexpected outcomes, which will enable it to make more accurate predictions and, therefore, avoid surprises in the future. In our case, Figures 9 and 10 illustrate that the surprises associated with $BT$ (Barrier Tactile) and $D$ (Depth) are reduced in comparison to those observed prior to learning. This clearly indicates that the predictions made by the agent are more accurate with respect to the environment when the $HV$ factor is present. In summary, before learning, the agent tends to go straight to the barrier and bumps into it in many occasions. Yet, after learning, the agent tends to avoid the barrier by side stepping around it in its trajectory to the target.
Figure 9: Surprise coefficients over BTBT (Barrier Tactile) before and after learning HVHV.
Figure 10: Surprise coefficients over $D$ (Depth) before and after learning HVHV.
Figure 11: Surprise coefficients over the Utility before and after learning HVHV.